ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 47, 46, 60 Peller, Vladimir Regularity conditions for vectorial stationary processes We discuss different regularity conditions for vectorial stationary processes with discrete time. We characterize such processes in terms of their spectral densities. We use the author’s results on the so-called recovery problem for unitary-valued matrix functions.